# GUing: A Mobile GUI Search Engine using a Vision-Language Model

### v2

@@ Abstract.
App developers use the Graphical User Interface (GUI) of other apps as a source of inspiration for designing and improving their own apps. Recent research has thus suggested retrieving relevant GUI designs that match a certain text query from screenshot datasets acquired through crowdsourced or automated exploration of GUIs. However, such text-to-GUI retrieval approaches only leverage the textual information of the GUI elements, neglecting visual information such as icons or background images. In addition, retrieved screenshots are not steered by app developers and often lack important app features that require particular input data.
To overcome these limitations, this paper proposes GUing, a GUI search engine based on a vision-language model called GUIClip, which we trained specifically for the problem of designing app GUIs. For this, we first collected from Google Play app introduction images which usually display the most representative screenshots and are often captioned (i.e. labeled) by app vendors. Then, we developed an automated pipeline to classify, crop, and extract the captions from these images. This resulted in a large dataset which we share with this paper: including 303k app screenshots, out of which 135k have captions. We used this dataset to train a novel vision-language model, which is, to the best of our knowledge, the first of its kind in GUI retrieval. We evaluated our approach on various datasets from related work and in manual experiment. The results demonstrate that our model outperforms previous approaches in text-to-GUI retrieval achieving a Recall@10 of up to 0.69 and a HIT@10 of 0.91. We also explored the performance of GUIClip for other GUI tasks including GUI classification and sketch-to-GUI retrieval with encouraging results.
@@ 1. Introduction
The Graphical User Interfaces (GUI) of mobile apps serve as the primary means of interaction between users and their devices. A well-designed GUI streamlines the navigation process, facilitates the accomplishment of user tasks, and improves the overall user experience— contributing towards a higher user engagement and retention (Hassan et al., 2020; Chen et al., 2021). Furthermore, a modern, attractive, and user-friendly GUI can potentially differentiate an app from its market counterparts, amplifying its likelihood of success in the highly competitive app market (Moran et al., 2018; Martens and Maalej, 2019b).
To design a good GUI, app developers may rely on multiple sources. In addition to initial requirements elicitation and prototyping, app developers often explore other GUIs of related apps. Numerous approaches have thus been suggested to retrieve relevant GUIs from existing apps. Researchers have proposed GUI retrieval approaches that accept sketches (Huang et al., 2019; Mohian and Csallner, 2022, 2023), wireframes (Deka et al., 2017; Liu et al., 2018; Chen et al., 2020), or screenshots (Li et al., 2021; Bunian et al., 2021) as input to locate similar or fitting designs. While certainly useful, these approaches require a preliminary graphical prototype, which might not be available or might be too restrictive in early ideation phases. Therefore, when only general ideas and textual requirements are available, text-to-GUI retrieval approaches would be particularly useful.
@@
Recent GUI search engines, such as GUIGLE (Bernal-Cardenas et al., 2019) and RaWi (Kolthoff et al., 2023), enable their users to search GUI datasets using text queries. Both engines use app metadata and GUI-related text for query matching. GUIGLE adopts basic keyword matching, while RaWi calculates semantic similarity based on a BERT model (Devlin et al., 2019). However, these text-based retrieval approaches solely use textual information available within the GUI such as text blocks or button labels, neglecting key visual information such as images, layouts, and backgrounds. Furthermore, the underlying GUI datasets (namely Rico (Deka et al., 2017) and ReDraw (Moran et al., 2020)) are created by crowdsourced or automated exploration of app screens at runtime. However, the access to certain screens may require authorization or initial configurations, which is often skipped in automated exploration. Even with a crowdsourced exploration, some app features may not be captured, particularly those requiring substantial or specific input data, such as dashboard pages with charts. Thus, screenshots crawled at runtime may fail to comprehensively capture essential features of the app.
In this paper, we propose a novel GUI search engine based on vision-language machine learning models trained with app introduction images from Google Play as shown on Figure 1. Recent Vision-Language Models (VLMs), such as CLIP (Radford et al., 2021), BLIP (Li et al., 2022a), and BLIP-2 (Li et al., 2023), are trained on large-scale image-caption data using contrastive learning. These models have the ability to transform images and text into a multimodal embedding, ensuring that semantically similar images and texts are mapped closely in the embedding space. Thus, by computing the similarity between the images and the textual query, these models can be used for text-to-image retrieval tasks. As prerequisite a large screenshot-caption dataset is needed. However, currently available datasets (Screen2Words (Wang et al., 2021) and Clarity (Moran et al., 2022)) are inadequate for training a VLM as our evaluation shows in Section 5. Therefore, we have created a new large screenshot-caption dataset.
@@
Google Play is one of the largest mobile app stores with thousands of apps from diverse domains. This makes it an rich source of inspiration for requirements elicitation and app design (Ferrari and Spoletini, 2023; Maalej et al., 2024). Particularly, the app introduction images on Google Play are a gold mine for GUI retrieval, as they are carefully selected and described by app developers to represent the important app features. Figure 2 shows examples of these images. A large portion can be categorized as surrounded screenshots, each displaying a screenshot and a succinct caption that describes the screen (Figure 2 (a)). In order to extract the screenshots and respective captions, we developed an automated pipeline, optimized for these images. The pipeline first classifies images into three categories as shown in Figure 2. Then it crops the screenshots area from the surrounded screenshots and extracts the caption from the surrounded screenshots (like the “LOG BLOOD PRESSURE” in Figure 2 (a)). We demonstrate the efficacy of our pipeline in Section 2.
By applying the pipeline on the introduction images of approximately 117k apps, we created a comprehensive dataset comprising 303k screenshots, of which 135k have captions. We refer to the collection of 303k screenshots as the ScreenRepo and the subset containing captions as SCapRepo. We then used SCapRepo together with the Screen2Words and Clarity datasets to fine-tune the CLIP model and create a VLM specific to the GUI domain. We call the new model GUIClip. Our evaluation results demonstrate a promising performance of GUIClip for text-to-image GUI retrieval, outperforming both text-only approaches and the CLIP model. Based on GUIClip, we developed GUing, a search engine, that accepts textual queries as input to retrieve relevant GUI images from our ScreenRepo dataset as well as the Rico and ReDraw datasets. Our engine can easily be extended with additional screen datasets. Our manual evaluation shows that the search engine recommends highly relevant GUIs. Additional experiments suggest that GUIClip is also beneficial for other GUI-related tasks, such as sketch-to-GUI retrieval and GUI classification. The paper provides the following contributions, with source code and datasets publicly available https://github.com/Jl-wei/guing:
@@ 2. Dataset Creation
Created two datasets: (1) Google Play Screenshot Caption (SCapRepo) including 135k screenshot-caption pairs for the training of our vision-language model and (2) Google Play Screenshot Repository (ScreenRepo) including 303k screenshots for the search engine repository. Figure 3 depicts the datasets creation pipeline. In the initial step, we collected app introduction images from Google Play. Subsequently, we developed an image classifier to categorize the images into screenshots, surrounded screenshots and irrelevant. The area in a surrounded screenshot containing a screenshots was precisely cropped from the surrounding by applying object detection. Additionally, we extracted the captions from the surrounded screenshots by applying Optical Character Recognition (OCR). The images classified as screenshots together with the cropped screenshots constitute the ScreenRepo, while the screenshot-caption pairs extracted from surrounded screenshots constitute the SCapRepo. The following describes each step in detail.
@@ 2.2. Image Classification
For subsequent analysis, we classified the images obtained from the prior phase into three categories as depicted in Figure 3. A subset of these images show the complete display of an app user interface: those are screenshots. In some images, only a segment represents a captured screenshot. Those are categorized as surrounded screenshots. The remaining images either do not show a screenshot or only a partial or a slanted screenshot. Those are classified as irrelevant. As a manual classification of around 400k images is unreasonable, we trained an image classifier to automatically classify these images into screenshots, surrounded screenshot, or irrelevant.
@@ 2.2.2. Creation of Training Data
To train and fine-tune the classification model, a labelled dataset is required. For this, we randomly sampled 3,615 images from the collected filtered dataset (see Section 2.1). To annotate these images, we created an annotation guide with examples to clarify the definition of screenshots, surrounded screenshot, and irrelevant. Two of the authors then annotated these images independently. For the entire labelled dataset, only 7 images had conflicting labels, primarily due to disagreements between the surrounded screenshot and irrelevant categories. These conflicts arose because the images in question did not contain a full screenshot but rather a UI component, such as a widget. After discussion, we reached a consensus to label images featuring a larger UI component (as thus potentially inspiring for a screen design) as surrounded screenshot, and the others as irrelevant.
@@ 2.2.3. Classification Performance
We evaluated the performance of the classifier using the labelled app introduction images from the previous step, which we split into a 80:20 ratio for training and testing, respectively. We trained the classification model using mini-batch gradient descent, with a batch size of 64 and AdamW optimiser with an initial learning rate of $2e^{-5}$. The classifier was trained in 5 epochs on a machine with a NVIDIA Tesla T4 GPU with 16 GB VRAM. In order to enhance the robustness of our results, we performed a 10-fold cross-validation by repetitively splitting the dataset into distinct training and testing subsets ten times in a randomized manner. We then computed the mean values of the precision, recall, and F1 score across these runs. The results are presented in Table 1 and show a top performance of our classifier, with an average F1 score of 0.964.
@@ 2.3. Screen Cropping
The surrounded screenshots are the images that contain screenshots with additional visual frames. To automatically crop the screenshot areas from these surrounded screenshots, we trained an object detection model. This model is capable of localizing the screenshot within the larger image and subsequently generating a bounding box. A bounding box is defined as a rectangular delineation that encases the area of interest: in this case, the screenshot. Given the bounding boxes inside the surrounded screenshots, we can easily crop the screenshot area using an image processing library.
@@ 2.3.1. Object Detection Model
We use DETR (DEtection TRansformer) (Carion et al., 2020) as screenshot detector. DETR is an end-to-end object detector, composed of a Convolutional Neural Network (CNN) backbone and a encoder-decoder transformer (Vaswani et al., 2017). Atop the transformer decoder, dual output heads are appended: a linear layer dedicated to categorize class labels and a multi-layer perceptron (MLP) tasked with the generation of bounding boxes for the object location. The model uses so-called object queries to detect objects in an image. Each object query is designed to search specifically for one particular object of interest in the given image. We set the number of queries to 1, as we are interested in the largest screenshot within a surrounded screenshot.
@@ 2.3.3. Detection Performance
Table 2. Accuracy of the screen localization (mean values for 10-fold cross-validation).
@@ 2.3.4. Applying the Detector
We trained a DETR model with all of the 1,768 labelled images, and applied this model to the 191,993 images categorized as surrounded screenshot. DETR predicted a bounding box that localize the screenshot for each image. We then cropped these images according to the bounding box. The screenshots cropped from the surrounded screenshot as well as the app introduction images previously classified as screenshots represent the ScreenRepo dataset.
@@ 2.4. Caption Recognition
The majority of the images, classified as surrounded screenshots, include captions that provide a succinct descriptions of the screenshots. Typically, these captions are positioned either above or below the image. In order to accurately extract this textual information, we employed Optical Character Recognition (OCR), a commonly used technology for recognizing text displayed within images.
@@ 3.1.2. Model Architecture
GUIClip consists of two components, an image encoder and a text encoder. The image encoder, which adopts a Vision Transformer (ViT) model (Dosovitskiy et al., 2021), ingests image data and generates its corresponding image embedding. The text encoder is a transformer encoder (Vaswani et al., 2017) which accepts text as input to produce an associated text embedding. Employing contrastive learning techniques, the image and text embeddings are mapped into a multi-modal embedding space. In this shared space, images and texts presenting semantic similarity are mapped close to each other.
Contrastive learning on image-text pairs constructs a bridge between visual perception and natural language. Given a batch of N (image, text) pairs, the training objective is to identify the authentic pairings from $N\times N$ potential (image, text) combinations within the batch (Radford et al., 2021). To achieve this objective, the model formulates a multi-modal embedding space, concurrently training an image encoder and text encoder to maximise the cosine similarity of the image and text embeddings derived from the N authentic pairs in the batch. Simultaneously, it minimises the cosine similarity of the embeddings from the remaining $N^{2}-N$ incorrect pairings.
Figure 5. Overview of contrastive learning to train our vision-language model GUIClip (adapted from (Radford et al., 2021)).
@@
Overview of contrastive learning to train our vision-language model GUIClip (adapted from (Radford et al., 2021)).

### v3

@@ Abstract.
Graphical User Interfaces (GUIs) are central to app development projects. App developers may use the GUIs of other apps as a means of requirements refinement and rapid prototyping or as a source of inspiration for designing and improving their own apps. Recent research has thus suggested retrieving relevant GUI designs that match a certain text query from screenshot datasets acquired through crowdsourced or automated exploration of GUIs. However, such text-to-GUI retrieval approaches only leverage the textual information of the GUI elements, neglecting visual information such as icons or background images. In addition, retrieved screenshots are not steered by app developers and lack app features that require particular input data.
To overcome these limitations, this paper proposes GUing, a GUI search engine based on a vision-language model called GUIClip, which we trained specifically for the problem of designing app GUIs. For this, we first collected from Google Play app introduction images which display the most representative screenshots and are often captioned (i.e. labelled) by app vendors. Then, we developed an automated pipeline to classify, crop, and extract the captions from these images. This resulted in a large dataset which we share with this paper: including 303k app screenshots, out of which 135k have captions. We used this dataset to train a novel vision-language model, which is, to the best of our knowledge, the first of its kind for GUI retrieval. We evaluated our approach on various datasets from related work and in a manual experiment. The results demonstrate that our model outperforms previous approaches in text-to-GUI retrieval achieving a Recall@10 of up to 0.69 and a HIT@10 of 0.91. We also explored the performance of GUIClip for other GUI tasks including GUI classification and sketch-to-GUI retrieval with encouraging results.
Vision-Language Model, GUI Prototyping, Information Retrieval, Requirements Engineering
@@ 1. Introduction
The Graphical User Interfaces (GUI) of mobile apps serve as the primary means of interaction between users and their devices. A well-designed GUI streamlines the navigation process, facilitates the accomplishment of user tasks, and improves the overall user experience — contributing towards a higher user engagement and retention (Hassan et al., 2020; Chen et al., 2021). Furthermore, a modern, attractive, and user-friendly GUI can potentially differentiate an app from its market counterparts, amplifying its likelihood of success in the highly competitive app market (Moran et al., 2018; Martens and Maalej, 2019b). To design a good GUI, developers often create multiple prototypes with varying fidelity: from low-fidelity sketches for brainstorming to high-fidelity GUIs for testing and optimisation. GUI prototypes are used in interviews or workshops to discuss, refine, and validate requirements, helping reduce misunderstanding and ultimately saving resources.
In this context, app developers often explore other GUIs of related apps as a way for rapid prototyping and a source of inspiration to design and improve their own apps. Numerous approaches have thus been suggested to retrieve relevant GUIs from existing apps. Researchers have proposed GUI retrieval approaches that accept sketches (Huang et al., 2019; Mohian and Csallner, 2022, 2023), wireframes (Deka et al., 2017; Liu et al., 2018; Chen et al., 2020), or screenshots (Li et al., 2021; Bunian et al., 2021) as input to locate similar or fitting designs. While certainly useful, these approaches require a preliminary graphical prototype, which might not be available or might be too restrictive in early ideation phases. Therefore, when only general ideas and textual requirements are available, text-to-GUI retrieval approaches would be particularly useful.
@@
Recent GUI search engines, such as GUIGLE (Bernal-Cardenas et al., 2019) and RaWi (Kolthoff et al., 2023), enable their users to search GUI datasets using text queries. Both engines use app metadata and GUI-related text for query matching. GUIGLE adopts basic keyword matching, while RaWi calculates semantic similarity based on a BERT model (Devlin et al., 2019). However, these text-based retrieval approaches solely use textual information available within the GUI such as text blocks or button labels, neglecting key visual information such as images, layouts, and backgrounds. Furthermore, the underlying GUI datasets (namely Rico (Deka et al., 2017) and ReDraw (Moran et al., 2020)) are created by crowdsourced or automated exploration of app screens at runtime. However, the access to certain screens may require authorisation or initial configurations, which is often skipped in automated exploration. Even with a crowdsourced exploration, some app features may not be captured, particularly those requiring substantial or specific input data, such as dashboard pages with charts. Thus, screenshots crawled at runtime may fail to comprehensively capture essential features of the app.
In this paper, we propose a novel GUI search engine based on a vision-language model trained with app introduction images from Google Play as shown on Figure 1. Recent vision-language models, such as CLIP (Radford et al., 2021), BLIP (Li et al., 2022a), and BLIP-2 (Li et al., 2023), are trained on large-scale image-caption data using contrastive learning. These models have the ability to transform images and text into a multimodal embedding, ensuring that semantically similar images and texts are mapped closely in the embedding space. Thus, by computing the similarity between the images and the textual query, these models can be used for text-to-image retrieval tasks. As prerequisite a large screenshot-caption dataset is needed. Since, currently available datasets (Screen2Words (Wang et al., 2021) and Clarity (Moran et al., 2022)) are inadequate for training a vision-language model (as our evaluation shows in Section 5), we have created a new large screenshot-caption dataset.
@@
Google Play is one of the largest mobile app stores with thousands of apps from diverse domains. This makes it a rich source of inspiration for requirements elicitation and app design (Ferrari and Spoletini, 2023; Maalej et al., 2024). Particularly, the app introduction images on Google Play are a gold mine for GUI retrieval, as they are carefully selected and described by developers to represent the important app features. Figure 2 shows examples of these images. A large portion can be categorised as surrounded screenshots, each displaying a screenshot and a succinct caption that describes the screen (Figure 2 (a)). In order to extract the screenshots and respective captions, we developed an automated pipeline, optimised for these images. The pipeline first classifies images into three categories as shown in Figure 2. Then, from the surrounded screenshots it crops the screenshot areas and extracts the captions (like “LOG BLOOD PRESSURE” in Figure 2 (a)).
By applying the pipeline on the introduction images of approximately 117k apps, we created a comprehensive dataset comprising 303k screenshots, of which 135k have captions. We refer to the collection of 303k screenshots as the ScreenRepo and the subset containing captions as SCapRepo. We then used SCapRepo together with the Screen2Words and Clarity datasets to fine-tune the CLIP model and create a vision-language model specific to the GUI domain. We call the new model GUIClip. Our evaluation results demonstrate a promising performance of GUIClip for text-to-image GUI retrieval, outperforming both text-only approaches and the CLIP model. Based on GUIClip, we developed GUing, a search engine that accepts textual queries as input to retrieve relevant GUI images from our dataset as well as the Rico and ReDraw datasets. The engine can easily be extended with additional screenshots. Our manual evaluation shows that the search engine recommends highly relevant GUIs, which can serve as a valuable source of design inspiration and a tool for rapid prototyping and requirements refinement. Two additional experiments suggest that GUIClip is also beneficial for other GUI-related tasks, such as sketch-to-GUI retrieval and GUI classification. The paper provides the following contributions, with source code and datasets publicly available for research purposes https://github.com/Jl-wei/guing:
@@ 2. Dataset Creation
Created two datasets: (1) Google Play Screenshot Caption (SCapRepo) including 135k screenshot-caption pairs for the training of our vision-language model and (2) Google Play Screenshot Repository (ScreenRepo) including 303k screenshots for the search engine repository. Figure 3 depicts the datasets creation pipeline. In the initial step, we collected app introduction images from Google Play. Subsequently, we developed an image classifier to categorise the images into screenshots, surrounded screenshots and irrelevant. The area in a surrounded screenshot containing a screenshots was precisely cropped from the surrounding by applying object detection. Additionally, we extracted the captions from the surrounded screenshots by applying Optical Character Recognition (OCR). The images classified as screenshots together with the cropped screenshots constitute the ScreenRepo, while the screenshot-caption pairs extracted from surrounded screenshots constitute the SCapRepo. The following describes each step in detail.
@@ 2.2. Image Classification
For subsequent analysis, we classified the images obtained from the prior phase into three categories as depicted in Figure 3. A subset of these images show the complete display of an app user interface: those are screenshots. In some images, only a segment represents a captured screenshot. Those are categorised as surrounded screenshots. The remaining images either do not show a screenshot or only a partial or a slanted screenshot. Those are classified as irrelevant. As a manual classification of around 400k images is unreasonable, we trained an image classifier to automatically classify these images into screenshots, surrounded screenshot, or irrelevant.
@@ 2.2.2. Creation of Training Data
To train and fine-tune the classification model, a labelled dataset is required. For this, we initially sampled 5,000 random images from the collected filtered dataset (see Section 2.1). After eliminating duplicates, 3,615 unique images remained. To annotate these images, we created an annotation guide with examples to clarify the definition of screenshots, surrounded screenshot, and irrelevant. Two of the authors then annotated these images independently. For the entire labelled dataset, only 7 images had conflicting labels, primarily due to disagreements between the surrounded screenshot and irrelevant categories. These conflicts arose because the images in question did not contain a full screenshot but rather a UI component, such as a widget. After discussion, we reached a consensus to label images featuring a larger UI component (as thus potentially inspiring for a screen design) as surrounded screenshot, and the others as irrelevant.
@@ 2.2.3. Classification Performance
We evaluated the performance of the classifier using the labelled app introduction images from the previous step, which we split into a 80:20 ratio for training and testing, respectively. The goal of training a machine learning (ML) model is to minimise the loss function, which quantifies the difference between the model’s predicted output and the actual target values. We trained the classification model using mini-batch gradient descent, with a batch size [^1] Batch size refers to the number of data samples processed simultaneously during a single training step in ML. of 64, AdamW (Loshchilov and Hutter, 2019) optimiser [^2] Optimisers, like AdamW, are strategies used to minimise the loss function., and an initial learning rate [^3] Learning rate is a parameter in ML algorithms, controlling the training step size while moving toward a minimum of a loss function. of $2e^{-5}$. The classifier was trained in 5 epochs on a machine with a NVIDIA Tesla T4 GPU with 16 GB VRAM. To enhance the robustness of our results, we performed a 10-fold cross-validation by repetitively splitting the dataset into distinct training and testing subsets ten times in a randomised manner. We then computed the mean values of the precision, recall, and F1 score across all runs. The results presented in Table 1 show a top performance of our classifier, with an average F1 score of 0.964.
@@ 2.3. Screen Cropping
The surrounded screenshots are the images that contain screenshots with additional visual frames. To automatically crop the screenshot areas from these surrounded screenshots, we trained an object detection model. This model is capable of localising the screenshot within the larger image and subsequently generating a bounding box. A bounding box is defined as a rectangular delineation that encases the area of interest: in this case, the screenshot. Given the bounding boxes inside the surrounded screenshots, we can easily crop the screenshot area using an image processing library.
@@ 2.3.1. Object Detection Model
We use DETR (DEtection TRansformer) (Carion et al., 2020) as screenshot detector. DETR is an end-to-end object detector, composed of a Convolutional Neural Network (CNN) backbone and a encoder-decoder transformer (Vaswani et al., 2017). Atop the transformer decoder, dual output heads are appended: a linear layer dedicated to categorise class labels and a multi-layer perceptron (MLP) tasked with the generation of bounding boxes for the object location. The model uses so-called object queries to detect objects in an image. Each object query is designed to search specifically for one particular object of interest in the given image. We set the number of queries to 1, as we are interested in the largest screenshot within a surrounded screenshot.
@@ 2.3.3. Detection Performance
Table 2. Accuracy of the screen localisation (mean values for 10-fold cross-validation).
@@ 2.3.4. Applying the Detector
We trained a DETR model with all of the 1,768 labelled images, and applied this model to the 191,993 images categorised as surrounded screenshot. DETR predicted a bounding box that localise the screenshot for each mage. We then cropped these images according to the bounding box. The screenshots cropped from the surrounded screenshot as well as the app introduction images previously classified as screenshots represent the ScreenRepo dataset.
@@ 2.4. Caption Recognition
The majority of the images, classified as surrounded screenshots, include captions that provide a succinct descriptions of the screenshots. Typically, these captions are positioned either above or below the image. In order to accurately extract this textual information, we employed Optical Character Recognition (OCR), a commonly used technology for recognising text displayed within images.
@@ 3.1.2. Model Architecture
GUIClip consists of two components, an image encoder and a text encoder. The image encoder, which adopts a Vision Transformer (ViT) model (Dosovitskiy et al., 2021), ingests image data and generates its corresponding image embedding. The text encoder is a transformer encoder (Vaswani et al., 2017) which accepts text as input to produce an associated text embedding. Employing contrastive learning techniques, the image and text embeddings are mapped into a multimodal embedding space. In this shared space, images and texts presenting semantic similarity are mapped close to each other.
Contrastive learning on image-text pairs constructs a bridge between visual perception and natural language. Given a batch of N (image, text) pairs, the training objective is to identify the authentic pairings from $N\times N$ potential (image, text) combinations within the batch (Radford et al., 2021). To achieve this objective, the model formulates a multimodal embedding space, concurrently training an image encoder and text encoder to maximise the cosine similarity of the image and text embeddings derived from the N authentic pairs in the batch. Simultaneously, it minimises the cosine similarity of the embeddings from the remaining $N^{2}-N$ incorrect pairings.
Figure 5. Overview of contrastive learning to train our vision-language model GUIClip (adapted from Radford et al. (Radford et al., 2021)).
@@
Overview of contrastive learning to train our vision-language model GUIClip (adapted from Radford et al. (Radford et al., 2021)).